Minute of KAGRA commissioning meeting
- Date: 2019/5/28 15:00~16:00
- Attendance: Miyakawa, Tomura, Fabian, Oshino, Washima, Yokozawa, Nakano, Terrence, Dahan, Kozakai, Pan, Sawada, Nakano, Yamada
- Zoom: Miyoki, Takahashi, Eleonora, Inoue, Fujii
Summary of last week
- Cryo cooler for Yarm stopped and finesse was recovered.
- Xarm 1451(1) with ITMX 73 K ETMX 120 K
- Yarm 1312(5) with ITMY 65 K ETMY 75 K
- Cryo status and plan
- ETMs has no dew on 80K viewport but ITMs has still dew on their viewport.
- Results of finesse measurement looks a bit too high. Needs estimation of loss inside cavity for Yarm.
- We should keep finesse measurement for going back the cooling down ITMY.
- For DRMI preparations, optics, circuit, cabling, control models are almost ready.
- Type A control test should be done during DRMI.
- Type B and A needs some IFO as sensor.
- Write manual if you experienced something.
- CAL: MICH and DRMI calibration model is ready.
- Remote work should be managed by calling to control room or sending email to gokagra.
- Smooth transition for when watch dog is engaged has been implemented to Type B. It should be implemented to Type A.
When? -> Maybe tomorrow morning during maintenance time or the next week during DRMI.
- Engineering run:
- IMC is the first candidate which will be used for run. This is because Guardian is ready for only IMC, not for Michelson or X arm.
- Next meeting: 5/28.